A dc shunt motor after running for several hours on constant voltage mains of 400 V takes a field current of 1.6 A. If the temperature rise is known to be 40°C, what value of extra circuit resistance is required to adjust the field current to 1.6 A when starting from cold at 20°C? Temperature coefficient = 0.0043/°C at 20°C.

1
Expert's answer
2021-12-07T20:12:28-0500

resistance:

"R=R_{40}(1+\alpha(T-40))"


resistance for 40°C:

"R_{40}=400/1.6=250\ \Omega"


then resistance at 20°C:

"R=250(1+0.0043(20-40))=228.5\ \Omega"


so, value of extra circuit resistance:

"250-228.5=21.5\ \Omega"